# Multilingual Encoder Knows more than You Realize: Shared Weights Pretraining for Extremely Low-Resource Languages

### v1

@@ Abstract
Zeli Su1,2  Ziyin Zhang3  Guixian Xu1,2 $\dagger$  Jianing Liu2
[…]
@@ 1 Introduction
Figure 1: The relationship between population size and dataset size in OSCAR (y-axis, in MB) for various high-, middle-, and low-resource languages.
@@ 2.3 NLP for Minority Languages in China
To enhance the accessibility of minority languages in China, prior studies have primarily focused on curating annotated datasets for various NLP tasks. These efforts have mainly concentrated on three key task categories: text classification Qun et al. (2017); Sun et al. (2021a); Shi et al. (2023), question answering Sun et al. (2021b), and machine translation Zhang et al. (2024a). Prominent models specifically trained for these languages include CINO (Yang et al., 2022), MiLMo (Deng et al., 2023), and TiBert (Liu et al., 2022). However, despite such progress, none of these models have released their pre-training corpora, and there is still a notable gap in the availability of models capable of text generation in these languages.
@@ 3.2.2 Training Data
For machine translation, we leveraged Google Translate to create bidirectional translation pairs between Chinese and the minority languages (Tibetan, Uyghur, Kazakh, and Mongolian). These translations were verified by native speakers to ensure accuracy. A total of 2,000 sentence pairs from each language pair were selected to form the supplementary training data.
@@ 4.2.1 Experiment Setting
- MRC: We mainly use the TibetanQA dataset Sun et al. (2022) for this task, which claims to contain 20K examples. However, only 2K examples are publicly available. Thus we enrich it by integrating 5K examples from the TibetanSFT Corpus [^1] https://huggingface.co/datasets/shajiu/ParallelCorpusSFT and 3K examples translated from a Chinese MRC dataset Cui et al. (2019a) using Google Translate. This approach enables us to create a comprehensive dataset consisting of 10K examples.
@@
- MRC: For Chinese, we employ the CMRC 2018 dataset (Cui et al., 2019b), which consists of 10K samples. For Tibetan, we use 500 samples extracted from the publicly available TibetanQA dataset. For the other three minority languages (Uyghur, Mongolian, Kazakh), we utilize machine translation tools to translate and clean MRC data, ultimately selecting 500 samples per language.
@@
We employ two baseline models to ensure broad coverage and robust performance in handling Chinese minority languages. The first model builds upon LLaMA2-Chinese and is fine-tuned on the MC2 dataset, resulting in the MC2-LLaMA-13B model. The second baseline, referred to as mBART-CM, is an adaptation of mBART-cc25. Its vocabulary is expanded to include tokens specific to our minority languages, followed by further pretraining on MC2.
@@ 7 Limitations
Thus, we hope that future work will put more focus on the development of high-quality datasets in these minority languages and beyond, enabling a more thorough exploration of underrepresented languages in the LLM era.
@@ Appendix A Dataset Details
Language Data Size Number of Samples
Tibetan 2.2 GB 184,045
Uyghur 736 MB 90,441
Kazakh 937 MB 57,827
Mongolian 970 MB 171,847
Simplified Chinese 2.1 GB 836,075

### v2

@@ Abstract
[…]
https://github.com/asd765973346/xlm-swcm
https://huggingface.co/KEVVVV/xlm-swcm
@@
Zeli Su1,2thanks: Equal contribution.  Ziyin Zhang3*  Guixian Xu1,2thanks: Corresponding author.  Jianing Liu2
XU Han1,2  Ting Zhang1,2  Yushuang Dong1,2 1Key Laboratory of Ethnic Language Intelligent Analysis and Security Governance of MOE 2Minzu University of China  3Shanghai Jiao Tong University {rickamorty,guixian_xu,hanxu,jianing_liu,yushuangdong}@muc.edu.cn daenerystargaryen@sjtu.edu.cn      tozhangting@126.com
@@ 1 Introduction
Figure 1: The relationship between population size and dataset size in OSCAR (y-axis, in MB) for various high-, middle-, and low-resource languages. Kazakh (kk), Mongolian (mn), Tibetan (bo), and Uyghur (ug) represent the languages we studied in this work.
@@ 2.3 NLP for Minority Languages in China
To enhance the accessibility of minority languages in China, prior studies have primarily focused on curating annotated datasets for various NLP tasks. These efforts have mainly concentrated on three key task categories: text classification Qun et al. (2017); Shi et al. (2023), question answering Sun et al. (2021), and machine translation Zhang et al. (2024a). Prominent models specifically trained for these languages include CINO (Yang et al., 2022), MiLMo (Deng et al., 2023), and TiBert (Liu et al., 2022). However, despite such progress, none of these models have released their pre-training corpora, and there is still a notable gap in the availability of models capable of text generation in these languages.
@@ 3.2.2 Training Data
Machine Translation Data were constructed by leveraging Google Translate to create bidirectional translation pairs between Chinese and four minority languages (Tibetan, Uyghur, Kazakh, and Mongolian). Due to the scarcity of parallel corpora for these low-resource languages, we adopted different strategies: for Uyghur, Kazakh, and Mongolian, we manually aggregated fragmented datasets and converted existing English-minority language pairs to Chinese-minority pairs via reliable English-Chinese translation systems. For Tibetan, we sampled from the relatively abundant TibetanSFT Corpus.  [^1] https://huggingface.co/datasets/shajiu/ParallelCorpusSFTAll translations were verified by native speakers to ensure quality, with 2,000 sentence pairs selected per language to create balanced supplementary training data across all four minority languages.
@@ 4.2.1 Experiment Setting
- MRC: We mainly use the TibetanQA dataset Sun et al. (2022) for this task, which claims to contain 20K examples. However, only 2K examples are publicly available. Thus we enrich it by integrating 5K examples from the TibetanSFT Corpus and 3K examples translated from a Chinese MRC dataset Cui et al. (2019) using Google Translate. This approach enables us to create a comprehensive dataset consisting of 10K examples.
@@
- MRC: For Chinese, we employ the CMRC 2018 dataset (Cui et al., 2019), which consists of 10K samples. For Tibetan, we use 500 samples extracted from the publicly available TibetanQA dataset. For the other three minority languages (Uyghur, Mongolian, Kazakh), we utilize machine translation tools to translate and clean MRC data, ultimately selecting 500 samples per language.
@@
We employ two baseline models to ensure broad coverage and robust performance in handling Chinese minority languages. Both models are continually pretrained on the MC2 dataset. The first, MC2-LLaMA-13B, is based on LLaMA2-Chinese from previous work.Zhang et al. (2024b) The second, mBART-CM, is adapted from mBART-cc25 with an expanded vocabulary to include tokens specific to our minority languages.
@@ 6 Conclusion
Thus, we hope that future work will put more focus on the development of high-quality datasets in these minority languages and beyond, enabling a more thorough exploration of underrepresented languages in the LLM era. As more data becomes available and the model’s capabilities continue to improve, the exploration of these languages will become a key direction for future research.
Acknowledgements
This research was supported by the Joint Research Project of Li’an International Education Innovation Pilot Zone, Hainan Province, China (Grant No: 624LALH006).
@@ Appendix A Dataset Details
Language Data Size Number of Samples Tokens
Tibetan (bo) 2.2 GB 184,045 131,531,254
Uyghur (ug) 736 MB 90,441 85,514,016
Kazakh (kk) 937 MB 57,827 92,730,961
Mongolian (mn) 970 MB 171,847 142,370,017
Simplified Chinese (zh) 2.1 GB 836,075 519,615,707